OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2.

OpenStudy (radar):

y-(-3)=1/2(x-6) y+3=(1/2)(x-6)

OpenStudy (radar):

The point-slope form is (y1-y2)=m(x1-x2) Where m = slope.